find a pair of linear equations for parallel lines
2x-3y=5
2x-3y=-7

Answer by stanbon(75887)

find a pair of linear equations for parallel lines
2x-3y=5
2x-3y=-7
---------------------
The graphs of these two equations are parallel
because they have the same slope.
slope = 2/3 for both
=============================
Cheers,
Stan H.